OpenStudy (anonymous):

Whats the GCF for 25, 35 INewton?

OpenStudy (anonymous):

25 = 5 * 5 35 = 5 * 7 So the highest common factor is..

OpenStudy (anonymous):

easter egg

OpenStudy (anonymous):

175?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

uhh

OpenStudy (anonymous):

INewton I don't know...

OpenStudy (anonymous):

You seem to be gettign mixed up between HCF and LCM. HCF is highest Common Factor (in this case, 5). LCM is different

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Whats the lowest?

OpenStudy (anonymous):

Oh, 175! That's the answer right?

OpenStudy (anonymous):

easter egg

OpenStudy (anonymous):

Stop it!